one week she bought 1 pound of cashews and 2 pounds of pecans and paid $2.40
the next week she bought 4 pounds of almonds and one pound of cashews paying $3.60
a week later she bought 3 pounds of walnuts 1 pound of cashews and 1 pound of almonds for $2.10
how much would she have to pay on her next trip to the grocery store if she were to buy 1 pound of each 4 types of nuts? [hint: using the initial letter of the name oeach type of nut(i.e:c,p,a,and w)will ditinguish the unknown values in the equation that you form]

Answer by ankor@dixie-net.com(22740)

Let c = lbs of cashews
Let p = lbs of pecans
Let a = lbs of almonds
Let w = lbs of walnuts
:
Write an equation for each statement:
:
"she bought 1 pound of cashews and 2 pounds of pecans and paid $2.40"
1c + 2p = 2.40
:
"she bought 4 pounds of almonds and one pound of cashews paying $3.60"
4a + 1c = 3.60
:
"she bought 3 pounds of walnuts 1 pound of cashews and 1 pound of almonds for $2.10"
3w + 1c + 1a = 2.10
:
We have 3 equations with 4 unknowns, We have to make some assumptions:
All values have to be less than $1.00 Also assuming we are dealing with integer cents.
:
In the first equation, Assume that c and p = .80
.80 + 2(.80) = 2.40
:
Using c = .80 find a in the 2nd equation:
4a + .80 = 3.60
4a = 3.60 - .80
4a = 2.80
a = .70/lb for almonds
:
Using a = .70 and c = .80 find w:
3w + .70 + .80 = 2.10
3w = 2.10 - 1.50
3w = .60
w = .20/lb for walnuts
:
It seems to come out even, there may be other values that work.
:
how much would she have to pay on her next trip to the grocery store if she were to buy 1 pound of each 4 types of nuts?
.20 + .70 + .80 + .80 = $2.50